Yna Celestine believes that the average amount of time spent by her classmates




in studying their self learning module in Math per week is less than 300 minutes




with a standard deviation of 45 minutes. She took a random sample of 35 students




in their class and found out that average time spent for studying was 285 minutes.




Test the claim at the 0.05 level of significance.





1
Expert's answer
2022-05-24T23:09:43-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge300"

"H_1:\mu<300"

This corresponds to a one-tailed left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a left-tailed test is "z_c =- 1.6449."

The rejection region for this lrft-tailed test is "R = \{z:z<-1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{285-300}{45/\sqrt{35}}\approx-1.9720"

Since it is observed that "z=-1.9720<-1.6449=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(z<-1.9720)=0.024305," and since "p=0.024305<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 300, at the "\alpha = 0.05" significance level.